# https://byeongjun-park.github.io/Switch-DiT/

### v1

@@ Abstract
Diffusion models have achieved remarkable success across a range of generative tasks. Recent efforts to enhance diffusion model architectures have reimagined them as a form of multi-task learning, where each task corresponds to a denoising task at a specific noise level. While these efforts have focused on parameter isolation and task routing, they fall short of capturing detailed inter-task relationships and risk losing semantic information, respectively. In response, we introduce Switch Diffusion Transformer (Switch-DiT), which establishes inter-task relationships between conflicting tasks without compromising semantic information. To achieve this, we employ a sparse mixture-of-experts within each transformer block to utilize semantic information and facilitate handling conflicts in tasks through parameter isolation. Additionally, we propose a diffusion prior loss, encouraging similar tasks to share their denoising paths while isolating conflicting ones. Through these, each transformer block contains a shared expert across all tasks, where the common and task-specific denoising paths enable the diffusion model to construct its beneficial way of synergizing denoising tasks. Extensive experiments validate the effectiveness of our approach in improving both image quality and convergence rate, and further analysis demonstrates that Switch-DiT constructs tailored denoising paths across various generation scenarios.
@@ 5.1 Experimental Setup
We employed the AdamW optimizer [28] with a fixed learning rate of 1e-4 and no weight decay. Additionally, we applied random horizontal flips to the training data and used $\lambda_{dp}=1$. For the FFHQ dataset [22], we trained for 100k iterations, while for the ImageNet dataset [6], we trained for 400k iterations. All models were trained with a batch size of 256 on 8 NVIDIA A100 GPUs.
@@ 5.2 Switch-DiT Design
[…]
[Uncaptioned image] Figure 4: Correlation of GFLOPs and FID on ImageNet. Switch-DiT transcends the tradeoff of DiT.
@@ 5.3 Comparative Evaluation
Unconditional Image Generation on FFHQ [22] (Model Configuration: B)
DiT
DTR
Switch-DiT
Figure 6: Qualitative Comparison with the same configuration on FFHQ.
Class-conditional Image Generation on ImageNet [6]
S
B
XL
  Shetland Rotisserie Beagle Carbonara Alp Church Toaster
Figure 7: Qualitative results of Switch-DiT on ImageNet.
5.3.3 Qualitative results.
Figure 6 shows the qualitative comparison of DiT [31], DTR [30] and Switch-DiT. The results demonstrate that Switch-DiT generates more realistic images, further verifying the effectiveness of our proposed method.
@@ 5.4 Analysis
5.4.1 Qualitative analysis.
Figure 7 illustrates the impacts of scaling model size on sample quality. The results demonstrate that larger model sizes for Switch-DiT consistently produce better images across multiple classes.
5.4.2 Loss weight ablation.
[…]
0.1 7.42
1.0 7.12
10.0 7.58
Table 4: $\lambda_{dp}$ effects on FFHQ.
We investigate the impact of varying $\lambda_{dp}$. Table 4 presents the ablative results on the FFHQ dataset, and we observe that the performance peaks at $\lambda_{dp}=1$. Also, the performance differences are not significantly influenced by changes in $\lambda_{dp}$, suggesting that the diffusion prior loss provides robust supervision on the gating output.
@@
To further validate this characteristic, we conduct an experiment on the FFHQ dataset for Random Allocation – For each block, one expert is shared across all timesteps, while another expert is randomly assigned to the first $N$ columns of the activation map obtained from Eq. 7, and the last one is assigned to the complementary set of the second one. As shown in Table 5, the results demonstrate that leveraging both bipartite matching and diffusion prior loss proves more effective than random allocation. We provide more experiments in Appedix A and analysis in Appendix B.
@@ 0.A.2 Comparison for Model Configuration Large
Figure A: Denoising path across $\lambda_{dp}$ on FFHQ. We show the denoising paths of Switch-DiT-B for four timesteps ($0.25T,0.5T,0.75T,T$). We sort the expert index for simplicity, and experts in orange are common experts shared across all timesteps.
@@ 0.B.1 Scalability
Figure B shows the correlation between GFLOPs and the FID-50K score for the Switch-DiT. Similar to DiT [31], we also observe a strong correlation from Switch-DiT-S to Switch-DiT-L, confirming its scaling behavior. Notably, in terms of the FID score, Switch-DiT-L outperforms DiT-XL even in smaller model configurations (9.40 vs. 8.78), further demonstrating its scalability with the lower FID score of 8.76 achieved by Switch-DiT-XL.
0.B.2 Impacts on Loss Weight
Our gating networks are implicitly trained via the noise prediction loss ${\mathcal{L}}_{noise}$, and directly supervised by the diffusion prior loss ${\mathcal{L}}_{dp}$. Therefore, we investigate the impacts of the balancing term between these two losses (i.e., $\lambda_{dp}$) on the construction of the denoising path within the diffusion model.
Figure A shows the denoising paths of Switch-DiT-B on the FFHQ dataset. Interestingly, denoising paths for $\lambda_{dp}=0.1$ and $\lambda_{dp}=1$ are the same, while those for $\lambda_{dp}=10$ exhibits different paths. This discrepancy arises because the gating network does not fully reflect the training signal from ${\mathcal{L}}_{noise}$ when $\lambda_{dp}=10$. Instead, the determination of denoising paths is primarily influenced by the inherent randomness in the bipartite matching process. Additionally, we observe that the denoising paths converge within 500 iterations for $\lambda_{dp}=1$ and $\lambda_{dp}=10$, whereas it takes thousands of iterations for $\lambda_{dp}=0.1$. The slow convergence is attributed to the failure to promptly reflect inter-task relationships, thus leading to ineffective management of negative transfers between conflicting tasks during the early stages of learning. In contrast, when $\lambda_{dp}=1$, gating networks properly reflect training signals from noise prediction loss and inter-task relationships.
@@ 0.B.2 Impacts on Loss Weight
Figure C: Qualitative comparison on ImageNet. Switch-DiT-XL consistently better generates images across various class labels compared to DiT-XL [31] and DTR-XL [30].
0.B.3 Qualitative comparison
Figure C presents a qualitative comparison of the class-conditional image generation among DiT-XL [31], DTR-XL [30] and Switch-DiT-XL on the ImageNet dataset. From these results, we validate the effectiveness of Switch-DiT in generating better realistic images across various class labels.
0.B.4 Discussion
0.B.4.1 Ethics Statement
Our approach is one of the generative models, and it may carry significant societal implications, particularly in applications like deep fakes and addressing biased data. Moreover, training generative models on datasets containing biased or manipulated content inadvertently reinforce social biases, fostering the dissemination of deceptive information and shaping public perception. We support discussions in the research community to address these unintended consequences.
0.B.4.2 Reproducibility Statement
We have included our experimental code in the supplementary materials. We hope this assists in understanding our manuscript.
0.B.4.3 Limiatations and Future Works

### v2

# Switch Diffusion Transformer: Synergizing Denoising Tasks with Sparse Mixture-of-Experts

@@ Abstract
Diffusion models have achieved remarkable success across a range of generative tasks. Recent efforts to enhance diffusion model architectures have reimagined them as a form of multi-task learning, where each task corresponds to a denoising task at a specific noise level. While these efforts have focused on parameter isolation and task routing, they fall short of capturing detailed inter-task relationships and risk losing semantic information, respectively. In response, we introduce Switch Diffusion Transformer (Switch-DiT), which establishes inter-task relationships between conflicting tasks without compromising semantic information. To achieve this, we employ a sparse mixture-of-experts within each transformer block to utilize semantic information and facilitate handling conflicts in tasks through parameter isolation. Also, we propose a diffusion prior loss, encouraging similar tasks to share their denoising paths while isolating conflicting ones. Through these, each transformer block contains a shared expert across all tasks, where the common and task-specific denoising paths enable the diffusion model to construct its beneficial way of synergizing denoising tasks. Extensive experiments validate the effectiveness of our approach in improving both image quality and convergence rate, and further analysis demonstrates that Switch-DiT constructs tailored denoising paths across various generation scenarios. Our project page is available at https://byeongjun-park.github.io/Switch-DiT/.
@@ 5.1 Experimental Setup
We employed the AdamW optimizer [28] with a learning rate of 1e-4 and no weight decay. Also, we applied random horizontal flips and used $\lambda_{dp}=1$. We trained for 100k and 400k iterations for the FFHQ [22] and ImageNet [6] datasets, respectively. All models were trained with a batch size of 256 on 8 NVIDIA A100 GPUs.
@@ 5.2 Switch-DiT Design
[…]
DiT-S [31] 33M 6.06 44.28 32.31 0.41 0.53
DTR-S [30] 33M 6.06 37.43 38.97 0.47 0.54
Switch-DiT-S 36M 6.74 33.99 42.99 0.52 0.54
DiT-B [31] 131M 23.01 27.96 64.72 0.57 0.52
DTR-B [30] 131M 23.02 16.58 87.94 0.66 0.53
Switch-DiT-B 144M 26.55 16.21 88.14 0.68 0.53
DiT-XL [31] 675M 118.64 9.40 166.83 0.77 0.48
DTR-XL [30] 675M 118.69 10.85 158.45 0.73 0.47
Switch-DiT-XL 749M 132.72 8.76 169.17 0.79 0.48
Table 2: Comparison across model sizes on ImageNet. Our Switch-DiT achieves consistent performance improvements.
Figure 4: Correlation of GFLOPs and FID on ImageNet. Switch-DiT transcends the tradeoff of DiT.
@@ 5.4 Analysis
To further validate this characteristic, we conduct an experiment on the FFHQ dataset for Random Allocation – For each block, one expert is shared across all timesteps, while another expert is randomly assigned to the first $N$ columns of the activation map obtained from Eq. 7, and the last one is assigned to the complementary set of the second one. As shown in Table 4, the results demonstrate that leveraging both bipartite matching and diffusion prior loss proves more effective than random allocation.
Please refer to the appendix for additional experiments and analyses on ablation studies concerning loss weight and qualitative comparisons.
@@ 6 Conclusion
Acknowledgements
This research was supported by Field-oriented Technology Development Project for Customs Administration through National Research Foundation of Korea (NRF) funded by the Ministry of Science & ICT and Korea Customs Service (NRF-2021M3I1A1097906).
@@ 0.A.2 Comparison for Model Configuration Large
[…]
DiT-S 92.07
DTR-S 42.24
Switch-DiT-S 34.39
Table C: Results on FFHQ.
0.A.3 High-resolution.
We further validate the effectiveness of our method on the FFHQ dataset with a higher resolution of 512 $\times$ 512. We trained DiT-S, DTR-S, and Switch-DiT-S for 100k iterations. As shown in Table C, our Switch-DiT achieves significant improvements compared to previous methods.
Class-conditional Image Generation on ImageNet [6]
S
B
XL
Shetland Rotisserie Beagle Carbonara Alp Church Toaster
Figure A: Qualitative results of Switch-DiT on ImageNet.
@@ 0.B.1 Scalability
Figure B shows the correlation between GFLOPs and the FID-50K score for the Switch-DiT. Similar to DiT [31], we also observe a strong correlation from Switch-DiT-S to Switch-DiT-L, confirming its scaling behavior. Notably, in terms of the FID score, Switch-DiT-L outperforms DiT-XL even in smaller model configurations (9.40 vs. 8.78), further demonstrating its scalability with the lower FID score of 8.76 achieved by Switch-DiT-XL. This implies that our SMoE-based architectural design is more effective for scaling diffusion models than simply increasing the hidden size and the number of transformer blocks.
Also, Fig. A illustrates the impacts of scaling model size on sample quality. The results demonstrate that larger model sizes for Switch-DiT consistently produce better images across multiple classes.
Figure C: Denoising path across $\lambda_{dp}$ on FFHQ. We show the denoising paths of Switch-DiT-B for four timesteps ($0.25T,0.5T,0.75T,T$). We sort the expert index for simplicity, and experts in orange are common experts shared across all timesteps.
Unconditional Image Generation on FFHQ [22] (Model Configuration: B)
DiT
DTR
Switch-DiT
@@
Figure D: Qualitative Comparison on FFHQ and ImageNet datasets.
0.B.2 Impacts on Loss Weight
[…]
0.1 7.42
1.0 7.12
10.0 7.58
Table D: $\lambda_{dp}$ effects on FFHQ.
We investigate the impact of varying $\lambda_{dp}$. Table D presents the ablative results on the FFHQ dataset, and we observe that the performance peaks at $\lambda_{dp}=1$. Also, the performance differences are not significantly influenced by changes in $\lambda_{dp}$, suggesting that the diffusion prior loss provides robust supervision on the gating output. However, we observe that the construction of the denoising path within the diffusion model changes as our gating networks are implicitly trained via the noise prediction loss ${\mathcal{L}}_{noise}$, and directly supervised by the diffusion prior loss ${\mathcal{L}}_{dp}$.
Figure C shows the denoising paths of Switch-DiT-B on the FFHQ dataset. Interestingly, denoising paths for $\lambda_{dp}=0.1$ and $\lambda_{dp}=1$ are the same, while those for $\lambda_{dp}=10$ exhibits different paths. This discrepancy arises because the gating network does not fully reflect the training signal from ${\mathcal{L}}_{noise}$ when $\lambda_{dp}=10$. Instead, the determination of denoising paths is primarily influenced by the inherent randomness in the bipartite matching process. Additionally, we observe that the denoising paths converge within 500 iterations for $\lambda_{dp}=1$ and $\lambda_{dp}=10$, whereas it takes thousands of iterations for $\lambda_{dp}=0.1$. The slow convergence is attributed to the failure to promptly reflect inter-task relationships, thus leading to ineffective management of negative transfers between conflicting tasks during the early stages of learning. In contrast, when $\lambda_{dp}=1$, gating networks properly reflect training signals from noise prediction loss and inter-task relationships.
0.B.3 Qualitative comparisons.
Figure D shows the qualitative comparison of DiT [31], DTR [30] and Switch-DiT. The results demonstrate that Switch-DiT generates more realistic images, further verifying the effectiveness of our proposed method.
Appendix 0.C Discussion
0.C.0.1 Ethics Statement
Our approach is one of the generative models, and it may carry significant societal implications, particularly in applications like deep fakes and addressing biased data.
0.C.0.2 Limiatations and Future Works